Calculate the temperature at which the root mean square velocity of a chlorine molecule will be equal to its escape velocity from the earth’s gravitational field. 

Given that the radius of earth is 6400 km.


1
Expert's answer
2021-04-13T06:42:21-0400

Solution.

"M(Cl)=35.5\sdot10^{-3}kgmol^{-1};"

"R_e=6.4\sdot 10^{6}m;"

"k=1.38\sdot10^{-23}JK^{-1};"

"N_A=6.02\sdot 10^{23}mol^{-1};"

"v-?;"

"W_k=\dfrac{mv^2}{2}; W_k=\dfrac{3kT}{2}; \implies" "v=\sqrt{\dfrac{3kT}{m}};"

"m=\dfrac{M}{N_A}; v=\sqrt{\dfrac{3kN_AT}{M}};"

"v_{escape}=\sqrt{2gR_e};"

"\sqrt{\dfrac{3kN_AT}{M}}=\sqrt{2gR_e} \implies T=\dfrac{2gR_eM}{3kN_{A}};"

"T=\dfrac{2\sdot 10mc^{-2}\sdot 6.4\sdot 10^6m\sdot 35.5\sdot 10^{-3}kgmol^{-1}}{3\sdot1.38\sdot10^{-23}JK^{-1}\sdot6.02\sdot 10^{23}mol^{-1}}=1.8\sdot10^5K;"


Answer: "T=1.8\sdot10^5K."